Find the moment of inertia of s pair of spheres, each having mass m and radius r, kept in contact about the tangent passing through the point of contact.plz explain the ans.

Grade:

1 Answers

Arun
5 years ago
Given in the question :-

Along the tangent ,The two bodies of mass m radius r are moving.
Refer to the attachment 
Here moment of inertia of one body about the axis(XY) passing by tangent.
 I = (2/5)mr²  

Now, Moment of inertia of first axis parallel to the other axis 
r = I + mr^2
r = (2/5) mr² + mr² 
r = (7/5)mr² .

Hence the moment of inertia of a pair of sphere is (7/5) mr
²